OpenStudy (anonymous):

Please help me... @Catch.me An alpha particle travelling at a velocity u makes an oblique elastic collision with a helium atom moving with a velocity which negligible compared to u.After collision , the alpha particle moves in the direction inclined at 30 degree to its initial direction of motion.What is the angle between the directions of motion of the helium atom and the initial direction of motion of alpha particle?

OpenStudy (anonymous):

Do you know dimension collision then just follow this instructions here http://www.physicsforums.com/showthread.php?t=415537

OpenStudy (anonymous):

Not really understand dimension collision...

OpenStudy (anonymous):

momentum is a vector apply conservation of momentum for x direction and y direction.

OpenStudy (anonymous):

 get their component and you will get the unknown theta. x direction mu = mv1cos 30 + Mv2cos theta . y direction 0 =- Mv2sin theta +mv1sin 30 .

OpenStudy (anonymous):

And the mass of alpha particle and helium is known

OpenStudy (anonymous):

I try first...

OpenStudy (anonymous):

I cant solve...

OpenStudy (anonymous):

@Catch.me

OpenStudy (anonymous):

Wait

OpenStudy (anonymous):

oh uh....alpha particle and helium has the same mass right.

OpenStudy (anonymous):

Then the angle angles after collision should be 90 degree thus the angle is 60 degree.

OpenStudy (anonymous):

@MichelleShum you don't know that right??

OpenStudy (anonymous):

yes...i don't know...

OpenStudy (anonymous):

@MichelleShum OK anyway the answer is 60 NOW we are going to prove that the are 90 degree.

OpenStudy (anonymous):

k...

OpenStudy (anonymous):

we got that conservation of kinetic energy, \[ m u = m V1 \cos 30 + mv2 \cos \theta.\] and \[0 = m v1 \sin 30 - m v2 \sin \theta. \] cancel masses in these equations and make mv2 in separated equation . \[V2 \cos \theta= u - v1 \cos 30 . \] \[V2 \sin \theta = v1 \cos 30.\] square both equations then add them sin ^2 + cos ^2 = 1 \[V _{2} ^{2} = u ^{2} - 2 u V _{1} \cos 30 + V _{1}^{2}.\] From kinetic equation \[0.5 u ^{2} = 0.5 V _{1}^{2} + 0.5 V _{2}^{2}\] now use v2^2 to substitute in kinetic equation. \[0 = 2 V _{1}^{2} - 2 V _{1} u \cos 30 \] \[V _{1} = u \cos 30 . \] \[u = \frac{ V _{1} }{ \cos 30 }\]Put that in your head we will use it later eq 1 . Now we will divide these equations we got them previous. \[V2 \sin \theta = v1 \cos 30.\] \[V2 \cos \theta= u - v1 \cos 30 . \] equal \[\tan \theta = \frac{ V _{1} \sin 30 }{ u - V _{1 \cos30} }\] use equation 1 instead of u \[\tan \theta = \frac{ \sin 30 }{ \frac{ 1 }{ \cos30 } - \cos30} = \frac{ \sin 30 }{ \frac{ 1- \cos ^{2} 30 }{ \cos 30 }} = \frac{ \sin30 }{ \frac{ \sin ^{2}30 }{ \cos 30 }} = \cot 30 \] Thus theta +30 = 90 degree WOW it's done.

OpenStudy (anonymous):

@MichelleShum Does it clear. i had to overcome some steps as they aren't tricky.

OpenStudy (anonymous):

ok...very clear...tq...

OpenStudy (anonymous):

Please help me this also... @Catch.me An aircraft is flying in a horizontal circle of radius 80km at a speed of 300m/s. What is the ratio of the lift on the aircraft to its weight?

OpenStudy (anonymous):

oh no this problem took me a lot of time and i have to go out now, so ask another one :D sorry.

OpenStudy (anonymous):

ok...nevermine...anyway thanks for your help...

OpenStudy (anonymous):

you are welcome. What is your course??

OpenStudy (anonymous):

I'm actually study Form 6...

OpenStudy (anonymous):

what is form 6 ??

OpenStudy (anonymous):

Something like A-level

OpenStudy (anonymous):

OK thanks @Mashy can you help here??

OpenStudy (anonymous):

yes i can help .. since the aircraft is taking the circle.. it must be banked right?